OpenStudy (anonymous):

1. Fill in the blank. 2n>14 n>_____ 2. n/5 ≥ 11 n≥ ____ 3. -3y ≤ -18 y≥ _____

OpenStudy (here_to_help15):

Hey :P

OpenStudy (here_to_help15):

n>14/2 n>7

OpenStudy (here_to_help15):

n/5>=11 n>=11*5 n>=55

OpenStudy (here_to_help15):

-3y<=-18 0<=-18+3y 18<=3y 18/3<=y 6<=y y>=6

OpenStudy (anonymous):

thanks

OpenStudy (here_to_help15):

np :)

OpenStudy (here_to_help15):

and i said Hey :P

OpenStudy (here_to_help15):

im starving

OpenStudy (here_to_help15):

ima go get little ceaser bb

OpenStudy (anonymous):

okay..

OpenStudy (here_to_help15):

Wanna slice :P